# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
Recall that throughout this work, the functions of interest belong to the space $CDF(\overline{\mathbb{R}})$ of CDF’s of extended-real-valued random variables. (Recall that this is nothing other than the space of nondecreasing, right-continuous functions taking values in $[0,1]$.) This section covers some specific simplifications of viscosity solutions theory that are applicable as long as the sub- or supersolutions in question are nondecreasing in the spatial variable. Specifically, it shows that, in this case, there is no loss of generality assuming that the test functions involved are themselves nondecreasing in space and defined globally in space.
@@
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Fix $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and let $\varphi$ be a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})>0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi_{1}(x)=\psi_{1}(x_{0}-r)$ for each $x\leq x_{0}-r$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:\mathbb{R}\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
@@
and equality holds if and only if $(x,t)=(x_{0},t_{0})$.
@@
By Taylor expansion, there is an $r>0$ and degree two polynomials $R$ and $Q$ such that $Q(0)=0$ and
$F(x,t)\leq\varphi(x,t)\leq R(x)+Q(t-t_{0})\quad\text{for each}\,\,(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$
Adding a term of the form $(x-x_{0})^{2}+(t-t_{0})^{2}$ to $R+Q$, there is no loss of generality in assuming in addition that $\varphi(x,t)=R(x)+Q(t-t_{0})$ if and only if $(x,t)=(x_{0},t_{0})$. Since $0<\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}R(x_{0})$, after possibly making $r$ smaller, it can similarly be assumed that $\partial_{x}R(x)>0$ for each $x\in[x_{0}-2r,x_{0}+2r]$.
Given $a>0$, let $\rho:\mathbb{R}\to\mathbb{R}$ be a smooth function such that $\rho(x)=0$ if $|x-x_{0}|\leq\frac{a}{2}$ and $\rho(x)=1$ if $|x-x_{0}|\geq a$. Define $\psi_{1}$ by
$\psi_{1}(x)=(1-\rho(x))R(x)+A\rho(x)\mathbf{1}_{[x_{0},+\infty)}(x)+R(x_{0}-\frac{a}{2})\rho(x)\mathbf{1}_{(-\infty,x_{0}]}(x).$
Since $R(x_{0})=\varphi(x_{0},t_{0})<A$, $\psi_{1}$ is a nondecreasing function provided $a$ is small enough. Fix such a small $a$ henceforth and let $\psi_{2}(t)=Q(t-t_{0})$. Since $F$ is nondecreasing in the spatial variable and $F(x,t)\leq\varphi(x,t)$ for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, the function $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)$ has the desired properties. ∎
@@
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Assume that $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and $\varphi$ is a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi(x)=\psi(x_{0})$ for each $x\leq x_{0}$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:[t_{0}-r,t_{0}+r]\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
@@
then
@@
$\varphi(x,t)$ $=\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{1}{3!}\partial^{3}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{3}$
$\quad+\frac{1}{4!}\partial^{4}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{4}+o(|x-x_{0}|^{4}+|t-t_{0}|^{4}),$
where $Q$ is some polynomial of degree four satisfying $Q(0)=0$. Since $F$ is nondecreasing in the $x$ variable; $F(x_{0},t_{0})=\varphi(x_{0},t_{0})$; and $F(x,t)\leq\varphi(x,t)$ for $(x,t)$ close enough to $(x_{0},t_{0})$, it follows that
@@
$\varphi(x,t)$ $\leq\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4}$
@@
$\psi(x,t)=h(x)+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}.$
Since $F\leq\varphi\leq P$ in $x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, $F$ is nondecreasing in the $x$ variable, and $F<A$ globally, it follows that $F\leq\psi$ in $\mathbb{R}\times[t_{0}-r,t_{0}+r]$ with equality only if $x\leq x_{0}$ and $t=t_{0}$. Therefore, the desired conclusion follows with $\psi_{1}(x)=h(x)$ and $\psi_{2}(t)=Q(t-t_{0})$. ∎
@@ Appendix B Scaling Limits of Some Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
In the proof of the theorem, there is one technicality that needs to be overcome. When one checks that, for instance, the limit $F$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$, it is desirable to invoke the asymptotics of $\mathscr{L}_{\epsilon}\varphi$ for an arbitrary smooth test function $\varphi$ that touches $F$ from above. If $\partial_{x}\varphi>0$ at the contact point, then Lemma 2 above shows that $\varphi$ can be assumed to be nondecreasing in $\mathbb{R}$ and then $\mathscr{L}_{\epsilon}\varphi$ is well-defined. Otherwise, if $\partial_{x}\varphi=0$ at the contact point, of course, $\varphi$ generally will not be monotone at all, and some work is required to proceed.
At the same time, smooth nondecreasing functions have a particular property that is suggestive here: If $\Phi$ is a smooth nondecreasing function in $\mathbb{R}$, and if $\partial_{x}\Phi(x_{0})=0$ at some point $x_{0}$, then $\partial_{x}^{2}\Phi(x_{0})=0$. This motivates the following lemma (which applies even if the subsolution is not monotone in the spatial variable).
@@
Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument and $Q:\mathbb{R}\to\mathbb{R}$ be continuous. An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
@@
Now Lemma 4 can be leveraged in conjunction with Lemma 3 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then the lemma shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
@@ B.1. Proof of Theorem 3
It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\frac{y}{\delta^{(N)}}-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
@@
Throughout what follows, define functions $c_{-},c_{+}:[0,\infty)\to[0,1]$ by
(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{+}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
Assumption (vi) ensures that these ODE’s are wellposed. For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
@@
(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,|N^{-1}n-t|+N^{-1}\leq\delta,\,\,t\in[0,T]\right\}=0.$
@@
Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v),
$N\mathscr{L}^{(N)}F_{0}^{(N)}=N\mathscr{L}^{(N)}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R^{(N)}$
@@
Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}N^{-1}(1+N^{-1}L)^{n}$ uniformly in $(x,n)$ as $N\to+\infty$. In view of the definition of $F_{N}$, this implies that, for each $T>0$,
$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}N^{-1}(1+LN^{-1})^{[NT]+1}.$
@@
Step 3: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that $F^{\star}$ coincides with $F_{\text{in}}$ in the sense of (65). Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
@@
Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain. The goal is to show that
@@
Step 3, Case (I). First, consider case (I). Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{*}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$. In particular, $x_{0}$ is a local minimum of $x\mapsto\varphi(x,t_{0})$, hence
@@
if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$. (That $Q(0)\leq 0$ is a consequence of the fact that the function $c\mapsto c-Q(c)$ maps $[0,1]$ into itself by assumptions (i) and (iv); see the discussion in Section 5.1.) Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
@@
Therefore, assume henceforth that (b) holds. According to Lemma 3 with $A=1-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that
$\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
$G^{\star}(x,t)\leq\varphi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
$(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad 1-\delta<c_{*}\leq\varphi(+\infty,t)\leq 1-\frac{\delta}{4}.$

### v3

@@ A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
Recall that throughout this work, the functions of interest belong to the space $CDF(\overline{\mathbb{R}})$ of CDF’s of extended real-valued random variables. (Recall that this is nothing other than the space of nondecreasing, right-continuous functions taking values in $[0,1]$.) This section covers some specific simplifications of viscosity solutions theory that are applicable as long as the sub- or supersolutions in question are nondecreasing in the spatial variable. Specifically, it shows that, in this case, there is no loss of generality assuming that the test functions involved are both globally defined and nondecreasing in the spatial variable.
@@
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Fix $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ and let $\varphi$ be a smooth function defined in some neighborhood of $(x_{0},t_{0})$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})>0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi_{1}(x)=\psi_{1}(x_{0}-r)$ for each $x\leq x_{0}-r$ and $\psi_{1}(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:\mathbb{R}\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
@@
and equality holds if and only if $(x,t)=(x_{0},t_{0})$. Further, for any $\delta>0$, it is possible to construct $\psi$ in such a way that
$\psi(x_{0},t_{0})=\varphi(x_{0},t_{0}),\quad\partial_{x}\psi(x_{0},t_{0})=\partial_{x}\varphi(x_{0},t_{0}),\quad\partial_{t}\varphi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0}),$
$\left|\partial_{x}^{2}\psi(x_{0},t_{0})-\partial_{x}^{2}\varphi(x_{0},t_{0})\right|\leq\delta.$
@@
Fix $\delta>0$. By Taylor’s theorem, there is an $r>0$ and second-degree polynomials $R$ and $S$ such that $\partial^{m}_{x}R(x_{0})=\partial_{x}^{m}\varphi(x_{0},t_{0})$ for $m\in\{0,1\}$, $S(0)=0$, $\partial_{t}S(0)=\partial_{t}\varphi(x_{0},t_{0})$, and
$F(x,t)\leq\varphi(x,t)\leq R(x)+S(t-t_{0})\quad\text{for each}\,\,(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r].$
Further, $R$ can be chosen in such a way that $|\partial_{x}^{2}R(x_{0})-\partial_{x}^{2}\varphi(x_{0},t_{0})|\leq\delta/2$. Adding a term of the form $\delta(x-x_{0})^{2}/4+t^{2}$ to $R+S$, there is no loss of generality in assuming in addition that $\varphi(x,t)=R(x)+S(t-t_{0})$ if and only if $(x,t)=(x_{0},t_{0})$. Since $0<\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}R(x_{0})$, after possibly making $r$ smaller, it can similarly be assumed that $\partial_{x}R(x)>0$ for each $x\in[x_{0}-2r,x_{0}+2r]$.
Given $a\in(0,r)$, let $\rho:\mathbb{R}\to\mathbb{R}$ be a smooth function such that $\rho(x)=0$ if $|x-x_{0}|\leq\frac{a}{2}$ and $\rho(x)=1$ if $|x-x_{0}|\geq a$. Define $\psi_{1}$ by
$\psi_{1}(x)=(1-\rho(x))R(x)+A\rho(x)\mathbf{1}_{[x_{0},+\infty)}(x)+R\left(x_{0}-\frac{a}{2}\right)\rho(x)\mathbf{1}_{(-\infty,x_{0}]}(x).$
Since $R(x_{0})=\varphi(x_{0},t_{0})<A$, $\psi_{1}$ is a nondecreasing function provided $a$ is small enough. Fix such a small $a$ henceforth and let $\psi_{2}(t)=S(t-t_{0})$. Since $F$ is nondecreasing in the spatial variable and $F(x,t)\leq\varphi(x,t)$ for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, the function $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)$ has the desired properties. ∎
@@
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Assume that $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ and $\varphi$ is a smooth function defined in some neighborhood of $(x_{0},t_{0})$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi(x)=\psi(x_{0})$ for each $x\leq x_{0}$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:[t_{0}-r,t_{0}+r]\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
@@
then $\psi(x_{0},t_{0})=\varphi(x_{0},t_{0})$, $\partial_{t}\psi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0})$,
@@
$\varphi(x,t)$ $=\varphi(x_{0},t_{0})+S(t-t_{0})+\frac{1}{3!}\partial^{3}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{3}$
$\qquad+\frac{1}{4!}\partial^{4}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{4}+o(|x-x_{0}|^{4}+|t-t_{0}|^{4}),$
where $S$ is some polynomial of degree four satisfying $S(0)=0$. Since $F$ is nondecreasing in the $x$ variable; $F(x_{0},t_{0})=\varphi(x_{0},t_{0})$; and $F(x,t)\leq\varphi(x,t)$ for $(x,t)$ close enough to $(x_{0},t_{0})$, it follows that
@@
$\varphi(x,t)$ $\leq\varphi(x_{0},t_{0})+S(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4}$
@@
$\psi(x,t)=h(x)+S(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}.$
Since $F\leq\varphi\leq P$ in $x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, $F$ is nondecreasing in the $x$ variable, and $F<A$ globally, it follows that $F\leq\psi$ in $\mathbb{R}\times[t_{0}-r,t_{0}+r]$ with equality only if $x\leq x_{0}$ and $t=t_{0}$. Therefore, the desired conclusion follows with $\psi_{1}(x)=h(x)$ and $\psi_{2}(t)=S(t-t_{0})+(t-t_{0})^{4}/4!$. ∎
@@ Appendix B Scaling Limits of Some Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
In the proof of the theorem, there is one technicality that needs to be overcome. When one checks that, for instance, the limit $F$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)-Q(F)$, it is desirable to invoke the asymptotics of $\mathscr{L}_{\epsilon}\varphi$ for an arbitrary smooth test function $\varphi$ that touches $F$ from above. If $\partial_{x}\varphi>0$ at the contact point, then Lemma 2 above shows that $\varphi$ can be assumed to be nondecreasing in $\mathbb{R}$ and then $\mathscr{L}_{\epsilon}\varphi$ is well-defined. Otherwise, if $\partial_{x}\varphi=0$ at the contact point, of course, $\varphi$ generally will not be monotone at all, and some work is required to proceed.
At the same time, smooth nondecreasing functions have a particular property that is suggestive here: If $G$ is a smooth nondecreasing function in $\mathbb{R}$, and if $\partial_{x}G(x_{0})=0$ at some point $x_{0}$, then $\partial_{x}^{2}G(x_{0})=0$. This motivates the following lemma (which applies even if the subsolution is not monotone in the spatial variable).
@@
Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument and $Q:\mathbb{R}\to\mathbb{R}$ be continuous. An upper semicontinuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
@@
Now Lemma 4 can be leveraged in conjunction with Lemma 3 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then Lemma 3 shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
@@ B.1. Proof of Theorem 3
It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,+\infty)$,
(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
@@
Throughout what follows, define functions $c_{-},c_{+}:[0,+\infty)\to\mathbb{R}$ by
(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{-}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
Assumption (vi) ensures that these ODE’s are wellposed. As discussed in Section 5.1, for any $N$, the function $q\mapsto q-Q^{(N)}(q)$ maps $[0,1]$ into itself. Taken together with assumption (vi), this implies that $Q(0)\leq 0$ and $Q(1)\geq 0$ for each $N$, and, thus,
$0\leq c_{+}(t),c_{-}(t)\leq 1\quad\text{for each}\,\,t\geq 0.$
For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
@@
(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,0\leq t\leq T,\,\,|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
@@
Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v), for each $y\in\mathbb{R}$,
$\mathscr{L}^{(N)}F_{0}^{(N)}(y)=\mathscr{L}^{(N)}F_{\text{in}}(y)=N^{-1}\mathcal{F}(\partial_{x}F_{\text{in}}(\delta^{(N)}y),\partial^{2}_{x}F_{\text{in}}(\delta^{(N)}y))+N^{-1}R^{(N)}(\delta^{(N)}y).$
@@
Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}L^{-1}\{(1+N^{-1}L)^{n}-1\}$ uniformly in $(x,n)$. In view of the definition of $F_{N}$, this implies that, for each $T>0$ and any $N\in\mathbb{N}$,
$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}L^{-1}(e^{LT}-1).$
@@
Step 3: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that
$\lim_{\delta\downarrow 0}\sup\left\{|F^{\star}(x,0)-F_{\text{in}}(y)|\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|x-y|\leq\delta\right\}=0,$
so $F^{\star}$ satisfies (65). Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$.
In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,+\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
@@
Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ in its domain. The goal is to show that
@@
Step 3, Case (I). First, consider case (I). Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{\star}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$. In particular, $x_{0}$ is a local minimum of $x\mapsto\varphi(x,t_{0})$, hence
@@
if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$. Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
@@
Therefore, assume henceforth that (b) holds. According to Lemma 3 with $A=1-\frac{\delta}{2}$, it is possible to fix an $r>0$ and a function $\psi:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ of the form
$\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
where $\psi_{1}$ is a nondecreasing $C^{2}$ function with $\psi_{1}(x)=\psi_{1}(x_{0})$ for $x\leq x_{0}$ and $\psi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\psi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\psi_{2}(t_{0})=0$; and
$\psi(x_{0},t_{0})=\varphi(x_{0},t_{0}),\quad\partial_{t}\psi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0}),$
$G^{\star}(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
$(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\psi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
Since $\psi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\psi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in(0,1)$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\psi(-\infty,t)\quad\text{and}\quad 1-\delta<c^{*}\leq\psi(+\infty,t)\leq 1-\frac{\delta}{4}.$